How many grams of I2 are present in a solution if 34.55 mL of 0.165 M Na2S2O3 solution is needed to titrate the I2solution?
1
Expert's answer
2018-09-27T05:47:57-0400
Formula:
CM=n/V,. n=CM x V
n= 0.03455 x 0,165
n= 0.00570075. Na2S2O3
nI= 0.00570075/2=0.002850375
m= 0.002850375 x 254= 0.72399525 g I2